OpenStudy (anonymous):

i have an algbrea quesionaolong with the answer can someone please tell me the solution to the these I have bothe the question and anser just need complete solution question(-5=6m-1) answer is -2-3 but how can you please break th is down for me

OpenStudy (turingtest):

please be careful int typing your question, it doesn't seem to make a lot of sense this way...

Parth (parthkohli):

Please rephrase that question.

OpenStudy (anonymous):

problem is -5=6m-1 I know that m=-2/3 but i need a breakdown of how they came up with this anwer

Parth (parthkohli):

Okay, gotcha! Have you heard of 'isolation'?

OpenStudy (anonymous):

no i am just trying to help a friend and ive been googling for help

Parth (parthkohli):

Well. You must get an answer like \(x = \text{a number}\). What you do is add, subtract, divide, multiply a number to both the sides in all steps such that you get the answer. Make sure that the number you use is not 0.

Parth (parthkohli):

For example, you have to solve\[x - 6 = 5 \]You will first add 6 to both sides\[x - 6 + 6 = 5 + 6 \\ x \cancel{-6 + 6}= 5 + 6\\ x = 11 \]

Parth (parthkohli):

And if you have\[2x + 11 = 3x + 12 \]You will first subtract 11 from both sides\[2x + 11 - 11 = 3x + 12 - 11\\2x = 3x+1 \\-1x = 1\\x = -1\]

OpenStudy (anonymous):

my hubby told me -1 was the answer to his problem but friend said its -2/3 is that the same thing

OpenStudy (anonymous):

hello are you still there

Parth (parthkohli):

Yes, it is -2/3. I gave you an example.

Parth (parthkohli):

What you do is: 1) add 1 to both sides. 2) divide both sides by 6.